OpenStudy (anonymous):

integral of 3 ((x^2)^1/3)

OpenStudy (anonymous):

Simplify the exponent, put the constant outside the integral, BAM easy integral.

OpenStudy (anonymous):

(x^2)^1/3 = x^ (2/3)

OpenStudy (anonymous):

@ncorrea92 Can you continue?

OpenStudy (callisto):

∫x^ndx = [x^(n+1) /(n+1)] +C Put n=2/3 into ^ Can you do it now?